OpenStudy (anonymous):

MEDALS AND FAN Determine whether the sequence converges or diverges. If it converges, give the limit. 48, 8, 4/3, 2/9 ... Converges; 288/5 Converges; 0 Diverges Converges; -12432

OpenStudy (anonymous):

it converges

OpenStudy (anonymous):

and 288/5 because a/(1-r) = 48/(1-1/6) = 48*6/5 =288/5

OpenStudy (anonymous):

it is an infinite G.P \[r=\frac{ 8 }{ 48 }=\frac{ 1 }{ 6 }<1\] \[s _{\infty }=\frac{ a }{ 1-r }\]